### v1

@@ 5.2 Evolution of scatter in the SHSMR
We now turn to the evolution of the scatter in the satellite SHSMR, which we explore in Figure 11. The first interesting fact from these figures is that when controlling for satellite history the scatter in the satellite SHSMR is significantly reduced. This is already apparent in Figure 10, in that $t_{\mathrm{infall}}$ clearly separates occupation in the $m_{\mathrm{sub}}-m_{\star}$ plane. The scatter in the present-day SHSMR decreases slightly for more recent infallers, and except for those galaxies with the longest preprocessing times it is independent of preprocessing time itself. This suggests, as does the discussion below, that it is the time spent as a satellite, $t_{\mathrm{sat}}$, the time that dominates the present-day SHSMR scatter. Whether the time spent as a satellite is spent within a less massive or more massive host is of secondary importance in these terms.
As the bottom panels show, when becoming satellites the SHSMR shows as little scatter as that of centrals. The middle panels, in turn, isolate the effect of preprocessing, which increases the scatter by up to 0.2 dex, with a timescale that depends on infall time (or, equivalently for this purpose, on $t_{\mathrm{sat}}$). That is, early infallers suffer preprocessing more quickly than late infallers, which seem to require several Gyr for the scatter to increase. This increase in scatter is, in the end, equally large as for recent infallers. As anticipated this suggests that $t_{\mathrm{sat}}$ is more important than preprocessing time in establishing both the mean and scatter of the SHSMR at infall.
As in the preceding section, we note that Figure 11 only shows satellites currently residing in $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$. Figure 16 shows the equivalent for satellites in $M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$ clusters. The only difference may be that the scatter in the SHSMR increases more quickly due to preprocessing for satellites currently residing in low-mass clusters, but the trends are otherwise consistent.
@@ 5.3 The evolution of segregation
Finally, we look at how the segregation described in Section 3.2 builds up over time. Figure 12 clearly shows the trend in Figure 3 evolving since infall, with galaxies both getting closer to the centre and losing preferentially more dark matter over time. Galaxies that fell most recently–within the last 2 Gyr or so—show little segregation within $R_{\mathrm{200m}}$, and cannot be found closer in than $0.1R_{\mathrm{200m}}$ as they have yet to reach the centre of the cluster. Except for the innermost radii, 10 Gyr after infall the SHSMR depends on cluster-centric distance as $\propto(R/R_{\mathrm{200m}})^{0.70}$. The flattening at $R<0.1R_{\mathrm{200m}}$ is due to our selection in present-day stellar mass, which means for heavily stripped subhaloes we only retain those with high $m_{\mathrm{sub}}/m_{\star}$. We confirm that selecting by infall mass instead produces relations consistent with power laws in the entire $R<R_{\mathrm{200m}}$ range, but the slopes at $R\gtrsim 0.1R_{\mathrm{200m}}$ do not change significantly. Using $t_{\mathrm{sat}}$ instead of $t_{\mathrm{infall}}$ (not shown) produces trends that match each other more closely than Figure 12, i.e., segregation does not depend as strongly on $t_{\mathrm{sat}}$ as it does on $t_{\mathrm{infall}}$. This is expected since preprocessing happens by definition while a galaxy is in the outskirts, or even entirely outside, of its (eventual) host cluster, well before galaxies have sunk into the cluster centre.
@@
Recent infallers show different subhalo-to-stellar mass ratios at fixed cluster-centric distance but this is no longer the case for $t_{\mathrm{infall}}\gtrsim 7$ Gya. Of such early infallers, indirect infallers have considerably lower ratios than direct infallers, due to the effect of the infalling host. Interestingly, the observation in Figure 3 that satellites have a lower $m_{\mathrm{sub}}/m_{\star}$ beyond $R_{\mathrm{200m}}$ is due to indirect infallers alone, while direct infallers have a constant $m_{\mathrm{sub}}/m_{\star}$ from $R_{\mathrm{200m}}$ outward. The galaxies that make it beyond $R_{\mathrm{200m}}$ after infall are those with the most elliptical orbits and therefore shortest pericentres, which suffer the most mass loss (e.g., Bakels et al., 2021; Smith et al., 2022). However, Haggar et al. (2023) have shown that galaxies in infalling groups are largely shielded from the host cluster if they are within the tidal radius of the infalling group, at least for some time. Combined, these effects explain why only indirect infallers show a decrease in $m_{\mathrm{sub}}/m_{\star}$ when located outside $R_{\mathrm{200m}}$ of the host cluster.
@@
We also note that the fact that earlier infallers tend to be located closer to the cluster centre was discussed by van den Bosch et al. (2016), who showed that the strong segregation of $m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{infall}}$ of dark matter subhaloes is due to the inside-out growth of haloes, i.e., that earlier infallers joined the host halo closer to the centre than late infallers do.
@@
Figure 12: Mass segregation build-up: subhalo-to-stellar mass ratio as a function of present-day 3-dimensional distance to the cluster centre, normalized by $R_{\mathrm{200m}}$ and split according to time since infall, as shown in the colourbar. Solid and dashed lines show direct and indirect infallers, respectively. The grey line shows a linear relation, for reference.
@@ 6 Quantifying mass loss in observations
Figure 13: Bias incurred in the estimation of mass loss by assuming it can be calculated as the ratio of satellite-to-central subhalo mass at fixed stellar mass (bottom). We calculate this bias as the ratio between the present-day-to-maximal subhalo mass (i.e., mass loss itself; top), and $\tau_{\mathrm{strip}}=1-m_{\mathrm{sat}}/m_{\mathrm{central}}$ (i.e., the observational proxy; middle). All curves are colour-coded by projected cluster-centric distance normalized by $R_{\mathrm{200m}}$, and shaded regions show corresponding normal uncertainties. Dashed lines correspond to using the outermost bin in the denominator instead; see Section 6 for details. Black lines show quantities for all satellites together.
@@
It is common among observational studies to attempt to quantify the mass loss in satellite galaxies by comparing their subhalo-to-stellar mass ratio with that of a sample of central galaxies with matching properties, usually stellar mass (e.g., Kumar et al., 2022). Under the assumption (justified, as we have seen) that stellar mass is much more stable than dark matter to mass loss, differences in the SHSMR between the two samples are taken as indication of mass loss. For instance, Wang et al. (2023) defined a stripping fraction as $\tau_{\mathrm{strip}}=1-m_{\mathrm{sub}}/m_{\mathrm{central}}$, i.e., assuming $m_{\mathrm{central}}=m_{\mathrm{sub}}^{\mathrm{max}}$ at fixed stellar mass, using the central SHSMR from Shan et al. (2017) to obtain $m_{\mathrm{central}}$. Alternatively, Niemiec et al. (2017) defined $\tau_{\mathrm{strip}}=1-m_{\mathrm{sub}}^{\mathrm{inner}}/m_{\mathrm{sub}}^{\mathrm{outer}}$, where $m_{\mathrm{sub}}^{\mathrm{inner}}$ and $m_{\mathrm{sub}}^{\mathrm{outer}}$ correspond to the mean masses of subhaloes in the inner and outer regions of clusters, respectively.
@@
As we have seen, stellar mass can increase quite significantly after infall. This means that the timescales for dark matter mass and stellar mass build-up in satellite galaxies are different, with maximal stellar mass attained typically 4 Gyr after maximal dark matter mass. Therefore, the points of comparison for dark matter loss and stellar mass loss are different. On the other hand, stellar mass generally contributes a small fraction of the total mass, even for heavily stripped subhaloes, so the stellar mass increase makes little difference for the overall amount of mass lost due to stripping.
@@
With the aim of providing an anchor for observational results such as the lensing measurements discussed in Section 1, here we evaluate how does the total, true mass loss compare to $\tau_{\mathrm{strip}}$. We calculate this mass loss bias as
$\mathrm{bias}=\frac{\tau_{\mathrm{strip}}}{1-m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}}-1=\frac{1-m_{\mathrm{sub}}/m_{\mathrm{central}}}{1-m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}}-1$ (2)
@@
where $m_{\mathrm{sub}}$ and $m_{\mathrm{central}}$ correspond to the subhalo masses of satellite and central galaxies, respectively (including baryons and dark matter). Figure 13 shows that this bias is approximately independent of stellar mass but strongly dependent on projected cluster-centric radius, increasing from 6% for the innermost bin shown to 51% in the outermost bin. [^11] These values correpsond to inverse-variance weighted averages of data points with $m_{\star}<5\times 10^{10}\,\mathrm{M}_{\odot}$, beyond which uncertainties dominate. In particular, note how the bias for $R_{01}>R_{\mathrm{200m}}$ diverges at higher stellar masses, because mass loss itself approaches zero. We also calculate $\tau_{\mathrm{strip}}$ similarly to Niemiec et al. (2017), by taking the ratios between mean subhalo masses in each of the three innermost radial bins and the outermost one, covering $0.8<R_{01}/R_{\mathrm{200m}}\lesssim 2$; these lead to the dashed lines. Because such a definition essentially ignores preprocessing, it leads to an underestimation of mass loss. Niemiec et al. (2017) used for the denominator a radial bin similar to the third radial bin in Figure 13, namely $1.00.55<R/(h^{-1}\,\mathrm{Mpc})<$ 1.0; because EAGLE galaxies in that radial bin have already lost 40–50% of their total mass already, this would lead to a larger underestimation of mass loss. We list the mean mass loss bias, Equation 2, in Table 3, when averaging each sample over the whole stellar mass range. Although the bias is significant it is well behaved in that it does not depend on stellar mass; it can therefore be treated as a correction to $\tau_{\mathrm{strip}}$ that depends only on the radial distribution of the satellites under consideration.
@@ 7 Conclusions
We study the mass content of satellite galaxies in galaxy clusters with masses $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ using an improved subhalo catalogue produced with the hbt+ algorithm (Han et al., 2018) applied to the 100 Mpc-box EAGLE cosmological hydrodynamical simulation, both at present and as a function of time. Consistent with previous work, we find satellites have a significantly reduced subhalo-to-stellar mass ratio compared to present-day central galaxies (Figure 1) and roughly 3 times higher intrinsic scatter in the subhalo-to-stellar mass relation at fixed stellar mass (Figure 4.) We find a strong dependence of the subhalo-to-stellar mass ratio on cluter-centric distance, independent of stellar mass, which is consistent with recent weak lensing measurements (Figure 3).
We also investigate the histories of present-day cluster galaxies. On average, satellite galaxies have spent roughly 5 Gyr within their current host. 30-40% of surviving satellites fell in to their present host indirectly as satellites of less massive systems, and this has a significant impact on their mass content: these indirect infallers had already lost 50–70% of their dark matter by infall, while direct infallers, i.e., those which fell into their current host as central galaxies, had lost $<10$%. Stellar mass, in contrast, continues to grow for roughly 2 Gyr after infall, and most satellites still have higher stellar masses today than they did at infall (Figures 7 and 8).
In fact, present-day satellites have lower SHSMR than centrals even at the time of maximal subhalo mass several Gya (Figure 10, lower panels), suggesting that while they were centrals they already experienced harsher environments than those galaxies which are still centrals today. Similarly, even galaxies in low-mass clusters (with $10^{13}<M_{\mathrm{200m}}/\mathrm{M}_{\odot}<5\times 10^{13}$) show significant signs of preprocessing, with the SHSMR at infall of indirect infallers being as low as 50% that of direct infallers into the same low-mass clusters at the same times (Figure 15). Overall, the effects of preprocessing—mass lost as a satellite of a previous host—and postprocessing—mass lost within the $z=0$ host—are equally important in shaping the SHSMR of present-day satellites.
@@
Finally, we quantify the degree to which mass loss can be estimated by comparing with the central SHSMR as done in weak lensing analyses. While this comparison provides a biased estimate, the bias is well behaved, with a steep dependence on cluster-centric radius but independent of stellar mass (Figure 13), suggesting it provides a good observational mass loss proxy.
@@ Acknowledgements
CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through FONDECYT Iniciación grant no. 11191125 and through BASAL project FB210003. JH is supported by NSFC (11973032, 11890691), 111 project (No. B20019), the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A03) and the Yangyang Development Fund.
@@
This work made use of Astropy: a community-developed core Python package and an ecosystem of tools and resources for astronomy (Robitaille et al., 2013; Collaboration et al., 2018, 2022, http://www.astropy.org). This work has benefited from open-source software including Matplotlib (Hunter, 2007, https://matplotlib.org/) NumPy (Harris et al., 2020, https://numpy.org/), SciPy (Virtanen et al., 2020, https://scipy.org/) Parts of the results in this work make use of the colormaps in the CMasher package (van der Velden, 2020, https://cmasher.readthedocs.io/index.html).

### v2

@@ 5.2 Evolution of scatter in the SHSMR
We now turn to the evolution of the scatter in the satellite SHSMR, which we explore in Figure 11. The first interesting fact from these figures is that when controlling for satellite history the scatter in the satellite SHSMR is significantly reduced. This is already apparent in Figure 10, in that $t_{\mathrm{infall}}$ clearly separates occupation in the $m_{\mathrm{sub}}-m_{\star}$ plane. The scatter in the present-day SHSMR decreases slightly for more recent infallers, and except for those galaxies with the longest pre-processing times it is independent of pre-processing time itself. This suggests, as does the discussion below, that it is the time spent as a satellite, $t_{\mathrm{sat}}$, the time that dominates the present-day SHSMR scatter. Whether the time spent as a satellite is spent within a less massive or more massive host is of secondary importance in these terms.
As the bottom panels show, when becoming satellites the SHSMR shows as little scatter as that of centrals. The middle panels, in turn, isolate the effect of pre-processing, which increases the scatter by up to 0.2 dex, with a timescale that depends on infall time (or, equivalently for this purpose, on $t_{\mathrm{sat}}$). That is, early infallers suffer pre-processing more quickly than late infallers, which seem to require several Gyr for the scatter to increase. This increase in scatter is, in the end, equally large as for recent infallers. As anticipated this suggests that $t_{\mathrm{sat}}$ is more important than pre-processing time in establishing both the mean and scatter of the SHSMR at infall.
As in the preceding section, we note that Figure 11 only shows satellites currently residing in $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$. Figure 16 shows the equivalent for satellites in $M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$ clusters. The only difference may be that the scatter in the SHSMR increases more quickly due to pre-processing for satellites currently residing in low-mass clusters, but the trends are otherwise consistent.
@@ 5.3 The evolution of segregation
Finally, we look at how the segregation described in Section 3.2 builds up over time. Figure 12 clearly shows the trend in Figure 3 evolving since infall, with galaxies both getting closer to the centre and losing preferentially more dark matter over time. Galaxies that fell most recently–within the last 2 Gyr or so—show little segregation within $R_{\mathrm{200m}}$, and cannot be found closer in than $0.1R_{\mathrm{200m}}$ as they have yet to reach the centre of the cluster. Except for the innermost radii, 10 Gyr after infall the SHSMR depends on cluster-centric distance as $\propto(R/R_{\mathrm{200m}})^{0.70}$. The flattening at $R<0.1R_{\mathrm{200m}}$ is due to our selection in present-day stellar mass, which means for heavily stripped subhaloes we only retain those with high $m_{\mathrm{sub}}/m_{\star}$. We confirm that selecting by infall mass instead produces relations consistent with power laws in the entire $R<R_{\mathrm{200m}}$ range, but the slopes at $R\gtrsim 0.1R_{\mathrm{200m}}$ do not change significantly. Using $t_{\mathrm{sat}}$ instead of $t_{\mathrm{infall}}$ (not shown) produces trends that match each other more closely than Figure 12, i.e. segregation does not depend as strongly on $t_{\mathrm{sat}}$ as it does on $t_{\mathrm{infall}}$. This is expected since pre-processing happens by definition while a galaxy is in the outskirts, or even entirely outside, of its (eventual) host cluster, well before galaxies have sunk into the cluster centre.
@@
Recent infallers show different subhalo-to-stellar mass ratios at fixed cluster-centric distance but this is no longer the case for $t_{\mathrm{infall}}\gtrsim 7$ Gya. Of such early infallers, indirect infallers have considerably lower ratios than direct infallers, due to the effect of the infalling host. Interestingly, the observation in Figure 3 that satellites have a lower $m_{\mathrm{sub}}/m_{\star}$ beyond $R_{\mathrm{200m}}$ is due to indirect infallers alone, while direct infallers have a constant $m_{\mathrm{sub}}/m_{\star}$ from $R_{\mathrm{200m}}$ outwards. The galaxies that make it beyond $R_{\mathrm{200m}}$ after infall are those with the most elliptical orbits and therefore shortest pericentres, which suffer the most mass loss (e.g. Bakels et al., 2021; Smith et al., 2022). However, Haggar et al. (2023) have shown that galaxies in infalling groups are largely shielded from the host cluster if they are within the tidal radius of the infalling group, at least for some time. Combined, these effects explain why only indirect infallers show a decrease in $m_{\mathrm{sub}}/m_{\star}$ when located outside $R_{\mathrm{200m}}$ of the host cluster.
@@
We also note that the fact that earlier infallers tend to be located closer to the cluster centre was discussed by van den Bosch et al. (2016), who showed that the strong segregation of $m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{infall}}$ of dark matter subhaloes is due to the inside-out growth of haloes, i.e. that earlier infallers joined the host halo closer to the centre than late infallers do.
@@
Figure 12: Mass segregation build-up. Subhalo-to-stellar mass ratio as a function of present-day three-dimensional distance to the cluster centre, normalized by $R_{\mathrm{200m}}$ and split according to time since infall, as shown in the colour bar. The solid and dashed lines show direct and indirect infallers, respectively. The grey line shows a linear relation, for reference.
@@ 6 Quantifying mass loss in observations
Figure 13: Bias incurred in the estimation of mass loss by assuming it can be calculated as the ratio of satellite to central subhalo mass at fixed stellar mass (bottom). This bias is calculated as the ratio of the present-day to maximum subhalo mass (i.e. mass loss itself; top), and $\tau_{\mathrm{strip}}=1-m_{\mathrm{sat}}/m_{\mathrm{central}}$ (i.e. the observational proxy; middle). All curves are colour-coded by projected cluster-centric distance normalized by $R_{\mathrm{200m}}$, and the shaded regions show corresponding normal uncertainties. The dashed lines correspond to using the outermost bin in the denominator instead; see Section 6 for details. The black lines show quantities for all satellites together.
@@
It is common among observational studies to attempt to quantify mass loss in satellite galaxies by comparing their subhalo-to-stellar mass ratio with that of a sample of central galaxies with matching properties, usually stellar mass (e.g. Kumar et al., 2022). Under the assumption (justified, as we have seen) that stellar mass is much more stable than dark matter to mass loss, differences in the SHSMR between the two samples are taken as indication of mass loss. For instance, Wang et al. (2023) defined a stripping fraction as $\tau_{\mathrm{strip}}=1-m_{\mathrm{sub}}/m_{\mathrm{central}}$, i.e. assuming $m_{\mathrm{central}}=m_{\mathrm{sub}}^{\mathrm{max}}$ at fixed stellar mass, using the central SHSMR from Shan et al. (2017) to obtain $m_{\mathrm{central}}$. Alternatively, Niemiec et al. (2017) defined $\tau_{\mathrm{strip}}=1-m_{\mathrm{sub}}^{\mathrm{inner}}/m_{\mathrm{sub}}^{\mathrm{outer}}$, where $m_{\mathrm{sub}}^{\mathrm{inner}}$ and $m_{\mathrm{sub}}^{\mathrm{outer}}$ correspond to the mean masses of subhaloes in the inner and outer regions of clusters, respectively.
@@
As we have seen, stellar mass can increase quite significantly after infall. This means that the timescales for dark matter mass and stellar mass build-up in satellite galaxies are different, with maximum stellar mass attained typically 4 Gyr after maximum dark matter mass. Therefore, the points of comparison for dark matter loss and stellar mass loss are different. On the other hand, stellar mass generally contributes a small fraction of the total mass, even for heavily stripped subhaloes, so the stellar mass increase makes little difference for the overall amount of mass lost due to stripping.
@@
With the aim of providing an anchor for observational results such as the lensing measurements discussed in Section 3.2, here we evaluate how the total true mass loss compares to $\tau_{\mathrm{strip}}$. We calculate this mass loss bias as
$\mathrm{bias}=\frac{\tau_{\mathrm{strip}}}{1-m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}}-1=\frac{1-m_{\mathrm{sub}}/m_{\mathrm{central}}}{1-m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}}-1,$ (2)
@@
where $m_{\mathrm{sub}}$ and $m_{\mathrm{central}}$ correspond to the subhalo masses of satellite and central galaxies, respectively (including baryons and dark matter). Figure 13 shows that this bias is approximately independent of stellar mass but strongly dependent on projected cluster-centric radius, increasing from 6% for the innermost bin shown to 51% in the outermost bin. [^12] These values correpsond to inverse-variance weighted averages of data points with $m_{\star}<5\times 10^{10}\,\mathrm{M}_{\odot}$, beyond which uncertainties dominate. In particular, note how the bias for $R_{xy}>R_{\mathrm{200m}}$ diverges at higher stellar masses, because mass loss itself approaches zero. We also calculate $\tau_{\mathrm{strip}}$ similarly to Niemiec et al. (2017), by taking the ratios between mean subhalo masses in each of the three innermost radial bins and the outermost one, covering $0.8<R_{xy}/R_{\mathrm{200m}}\lesssim 2$; these lead to the dashed lines. Because such a definition essentially ignores pre-processing, it leads to an underestimation of mass loss. Niemiec et al. (2017) used for the denominator a radial bin similar to the third radial bin in Figure 13, namely $1.00.55<R/(h^{-1}\,\mathrm{Mpc})<$ 1.0; because EAGLE galaxies in that radial bin have already lost 40–50% of their total mass already, this would lead to a larger underestimation of mass loss. We list the mean mass loss bias, Equation 2, in Table 3, when averaging each sample over the whole stellar mass range. Although the bias is significant it is well behaved in that it does not depend on stellar mass; it can therefore be treated as a correction to $\tau_{\mathrm{strip}}$ that depends only on the radial distribution of the satellites under consideration.
@@ 7 Conclusions
We studied the mass content of satellite galaxies in galaxy clusters with masses $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ using an improved subhalo catalogue produced with the hbt+ algorithm (Han et al., 2018) applied to the 100 Mpc-box EAGLE cosmological hydrodynamical simulation, both at present and as a function of time. Consistent with previous work, we find satellites have a significantly reduced subhalo-to-stellar mass ratio compared to present-day central galaxies (Figure 1) and roughly 3 times higher intrinsic scatter in the subhalo-to-stellar mass relation at fixed stellar mass (Figure 4). We find a strong dependence of the subhalo-to-stellar mass ratio on cluter-centric distance, independent of stellar mass, which is consistent with recent weak lensing measurements (Figure 3).
We also investigated the histories of present-day cluster galaxies. On average, satellite galaxies have spent roughly 5 Gyr within their current host. 30-40% of surviving satellites fell in to their present host indirectly as satellites of less massive systems, and this has a significant impact on their mass content: these indirect infallers had already lost 50–70% of their dark matter by infall, while direct infallers, i.e. those which fell into their current host as central galaxies, had lost $<10$%. Stellar mass, in contrast, continues to grow for roughly 2 Gyr after infall, and most satellites still have higher stellar masses today than they did at infall (Figures 7 and 8).
In fact, present-day satellites have lower SHSMR than centrals even at the time of maximum subhalo mass several Gya (Figure 10, lower panels), suggesting that while they were centrals they already experienced harsher environments than those galaxies which are still centrals today. Similarly, even galaxies in low-mass clusters (with $10^{13}<M_{\mathrm{200m}}/\mathrm{M}_{\odot}<5\times 10^{13}$) show significant signs of pre-processing, with the SHSMR at infall of indirect infallers being as low as 50% that of direct infallers into the same low-mass clusters at the same times (Figure 15). Overall, the effects of pre-processing (mass lost as a satellite of a previous host) and post-processing (mass lost within the $z=0$ host) are equally important in shaping the SHSMR of present-day satellites.
@@
Finally, we quantify the degree to which mass loss can be estimated observationally by comparing the satellite and central SHSMRs as done in weak lensing analyses. While this comparison provides a biased estimate, the bias is well behaved, with a steep dependence on cluster-centric radius but independent of stellar mass (Figure 13), suggesting it provides a good observational mass loss proxy.
@@ Acknowledgements
The code used for the analysis and to generate figures in this paper is publicly available at https://github.com/cristobal-sifon/eagle-satellites. The hbt+ catalogue is available upon request to J. Han.
We thank the anonymous referee for their useful suggestions. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through FONDECYT Iniciación grant no. 11191125 and through BASAL project FB210003. JH is supported by NSFC (11973032, 11890691), 111 project (No. B20019), the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A03) and the Yangyang Development Fund.
@@
This work made use of Astropy: a community-developed core Python package and an ecosystem of tools and resources for astronomy (Robitaille et al., 2013; Collaboration et al., 2018, 2022, http://www.astropy.org). This work has benefited from open-source software including Matplotlib (Hunter, 2007, https://matplotlib.org/) NumPy (Harris et al., 2020, https://numpy.org/), SciPy (Virtanen et al., 2020, https://scipy.org/), and Pandas (Reback2020pandas; Mckinney-proc-scipy-2010, https://pandas.pydata.org/). Parts of the results in this work make use of the colour maps in the CMasher package (van der Velden, 2020, https://cmasher.readthedocs.io/index.html).